OpenStudy (anonymous):

Convert f(x)=2x^2+2x+7 into vertex form.

OpenStudy (anonymous):

Complete the square to get vertex form

OpenStudy (anonymous):

A step-by-step would be fantastic

OpenStudy (anonymous):

Or a more detailed explanation...

OpenStudy (anonymous):

Alright, first factor out the number in front of x^2 2(x^2+x+7/2) Now take the coefficient of the x with no exponent, divide it by 2, and then square it. What do you get?

OpenStudy (anonymous):

.25?

OpenStudy (anonymous):

why are you dividing by 2 within that "2(x^2+x+7/2)?

OpenStudy (anonymous):

Yep, or 1/4. Now were gonna both add and subtract that after the x. 2(x^2+x+1/4-1/4+7/2)

OpenStudy (anonymous):

The coefficient in front of the x with no exponent, which is 1

OpenStudy (anonymous):

Its just what you do,

OpenStudy (anonymous):

\[\frac{2(x ^{2}+x+\frac{ 1 }{ 4 }-\frac{ 1 }{ 4 }+7}{2}\]

OpenStudy (anonymous):

with that () closed off. is that what it should look like?

OpenStudy (anonymous):

Oh no, the 2 is only under the 7

OpenStudy (anonymous):

Because first we factored out the 2

OpenStudy (anonymous):

so when we factor out a 2, we also divide 7 by 2?

OpenStudy (anonymous):

Yep we divided everything by 2 which is why the 2x^2+2x part became x^2+x

OpenStudy (anonymous):

I think i understand...

OpenStudy (anonymous):

so id be \[f(x) = 2 ( x ^{2} + x ) + \frac{ 7 }{ 2 }\]

OpenStudy (anonymous):

?

OpenStudy (anonymous):

Nope, how does the x^2+x+1/4 factor?

OpenStudy (anonymous):

In 2(x^2+x+1/4-1/4+7/2)

OpenStudy (anonymous):

(x+1/2)^2 right?

OpenStudy (anonymous):

So now we have 2((x+1/2)^2-1/4+7/2)

OpenStudy (anonymous):

-1/4+7/2=?